OpenStudy (anonymous):

Factor completely 81x2 - 4

OpenStudy (anonymous):

@Here_to_Help15

OpenStudy (alexandervonhumboldt2):

(9x-2)*(9x+2)

OpenStudy (anonymous):

how did yu get that?

OpenStudy (alexandervonhumboldt2):

a^2-b^2=(a+b)*(a-b)

OpenStudy (anonymous):

can you help with a few more?

OpenStudy (alexandervonhumboldt2):

9^2x^2-2^2=(9x+2)*(9x-2)

OpenStudy (alexandervonhumboldt2):

can you faan?

OpenStudy (here_to_help15):

3^4*x^2-2^2

OpenStudy (here_to_help15):

So yes he is right

OpenStudy (anonymous):

Which is a factor of 54xy + 45x + 18y + 15?

OpenStudy (anonymous):

and i already fanned both of you

OpenStudy (alexandervonhumboldt2):

dont know

OpenStudy (here_to_help15):

(2*3^3)xy + (3^2*5)x + (2*3^2)y+3.5 3(2.3^3*xy/3 + 3^2*5x/3+ 2*3^2*y/3+ 3.5/3 Answer- 3(6y+5)(3x+1)

OpenStudy (here_to_help15):

Answer: 3(6y+5)(3x+1)